# HBS - Hardware Build System: A Tcl-based, minimal common abstraction approach for build system for hardware designs

### v3

@@ 3.4. Ease of calling arbitrary programs in arbitrary places
The first use case originates from the complexity of designing hardware systems. To reduce manual work, engineers often automatically generate parts of the design. For example, register files are automatically generated from SystemRDL (Accellera SYSTEMS INITIATIVE, 2025). In such a case, there must be a way to call an external program generating hardware description of the register components. Existing indirect abstract hardware build systems allow calling external programs. However, they often require users to define a transitional so-called generator, which increases complexity and reduces legibility. Moreover, they usually do not allow to call external programs in arbitrary places.
The second use case includes support for EDA tools without the Tcl interface. Such tools are often open source simulators, for example, nvc (N., 2025), or verilator (W., 2025). These tools have interfaces similar to the standard software compiler. The user simply calls programs with proper arguments and proper order in a shell. Being able to call arbitrary programs in arbitrary places is required to support this kind of tools without an extra script or Makefile generation.
3.5. Independce from version control system
@@ 4. Hardware Build System architecture
This section describes the architecture of the implemented HBS. The system is available on GitHub (M., 2025) with the BSD 3-Clause License.
It was decided that HBS will be implemented using Tcl language because of the following reasons:
@@
    Implementing a hardware build system in Tcl allows the execution of build system code during the EDA tool flow. This, in turn, gives direct access to all EDA tool custom commands. Moreover, these custom commands can be evaluated in arbitrary places. This aspect satisfies requirements 3.1 and 3.4. The complexity of existing indirect abstract build systems for less common tasks results from executing build system code before EDA tool flow. The build system only prepares scripts that are later run by EDA tools.
@@
    If the EDA tool provides the Tcl interface, then the Tcl shell is provided by the EDA tool vendor. The shell is installed during the installation of vendor tools. This implies that, in some cases, the build system user does not even have to install additional programs. This aspect satisfies requirement 3.3.
@@
    Executing arbitrary programs in arbitrary places in Tcl is very simple. There is a dedicated exec command for invoking subprocesses. If executing a subprocess requires prior dynamic arguments evaluation, the exec command call must be prepended with the eval command. Even in Python, invoking a subprocess is not so straighforward. This Tcl feature satisfies requirement 3.4.
One of the most important things while designing the HBS was the separation of common build process operations that would constitute a common abstraction layer over EDA tools. At the end, it was decided that the following actions should constitute the common abstraction layer:
  1. (1)
    Target device setting - some EDA tools use the term “part” instead of “device,” for example, Vivado. Simulation EDA tools do not require a device setting. However, all synthesis EDA tools require information on the target device. This is why setting the device became part of the common abstraction layer.
  2. (2)
  3. (3)
    Library setting - setting HDL file library.
  4. (4)
    HDL standard setting - setting HDL file standard revision. The build system has to manage this because some tools can not handle analyzing different design units using different standard revisions, for example, nvc. In such a case, the build system must decide what the common standard revision shall be used for analyzing all HDL files.
  5. (5)
  6. (6)
  7. (7)
    Design top module setting - all EDA tools carrying out simulation or synthesis requires information on the top module.
  8. (8)
    Code generation - a hardware build system must provide a simple way for automatic arbitrary code generation. This is further explained in subsection 4.12.
4.1. General structure
The HBS is implemented in two files hbs.tcl and hbs. The first is implemented in Tcl, and the second in Python. The hbs.tcl file implements all the core features related directly to the interaction with the EDA tools. The hbs.tcl provides the following functions:
  1. (1)
    Dumping information about detected cores in JSON format.
  2. (2)
    Listing cores found in .hbs files.
  3. (3)
    Listing targets for a given core.
  4. (4)
    Running target provided as a command line argument.
The hbs is a wrapper for the hbs.tcl and serves the following additional functions:
  1. (1)
    Showing documentation for hbs Tcl symbols.
  2. (2)
    Generating dependency graph.
  3. (3)
    Listing testbench targets.
  4. (4)
    Running testbench targets.
By default, the user calls the hbs program. However, if none of the additional functions are required, the user can call the hbs.tcl directly. In such a case, the whole build system is limited to a single file. Check Table 1 to see how it compares to the size of existing hardware build systems.
@@ 4.2. Cores and cores detection
When the user executes hbs (or hbs.tcl), all directories, starting from the working directory, are recursively scanned to discover all files with the .hbs extension (symbolic links are also scanned). Files with the .hbs extension are regular Tcl files that are sourced by the hbs.tcl script. However, before sourcing .hbs files, the file list is sorted so that scripts with shorter path depth are sourced as the first ones. For example, let us assume the following three .hbs files were found:
@@
Such an approach allows controlling when custom symbols (Tcl variables and procedures) are ready to use. For example, if the user has a custom procedure used in multiple .hbs files, then the user can create separate utils.hbs file contaning utility procedures, and place it in the the project root directory. Within .hbs files, the user usually defines cores and targets, although the user is free to have any valid Tcl code in .hbs files.
@@
To register a core, the user must explicitly call hbs::Register procedure at the end of the core namespace. Such a mechanism helps to distinguish regular Tcl namespaces from Tcl namespaces representing core definitions. If the user forgets to register a core, the build system gives a potential hint. An example is presented in Listing 3.
Listing 3: An HBS error message with a hint on core registration.
[user@host tmp]\$ hbs run lib::core::tb
core lib::core not found, maybe the core is not registered (hsb::Register)
@@
    The user can easily stick to the VLNV identifiers if required. This is presented in Listing 4. In this case, the flip-flop core path is vendor::library::flip-flop::1.0.
@@
    The user can define arbitrary deep core paths (limited by the Tcl shell). This is presented in Listing 5. In this case, the core path consists of 7 parts.
@@
    The user can nest namespaces to imitate the structure of libraries and packages. This is presented in Listing 6. Three flip-flop cores are defined in the snippet. Listing 7 presents output for listing flip-flop cores.
Listing 4: Example flip-flop core definition with VLNV core path.
@@
Listing 5: Example flip-flop core definition with 7 core path parts.
@@
Listing 6: An example .hbs file with three flip-flops core definitions
@@
Listing 7: Listing flip-flop cores output.
[user@host tmp]\$ hbs list-cores
@@ 4.3. Targets and targets detection
HBS automatically detects targets. Targets are all Tcl procedures defined in the scope of core namespaces (namespaces with a call to hbs::Register). However, to allow users to define custom utility procedures within cores, procedures with names starting with the floor character (_) are not treated as core targets. Listing 8 presents an example edge detector core definition. The core path is vhdl::simple::edge-detector. The core has three targets: src, tb-sync, tb-comb, and one utility procedure _tb. The _tb procedure was defined to share calls common for testbench targets tb-sync and tb-comb. Moreover, all target procedures are also regular Tcl procedures. Such an approach allows for calling them in arbitrary places. The _tb procedure calls the src procedure because the edge_detector.vhd file is required for running testbench targets.
All targets are represented by a unique target path. Target path consists of the core path and target name. For example, the src target of the edge detector has the following path vhdl::simple::edge-detector::src.
Listing 8: Example edge detector core definition.
@@ 4.4. Testbench targets
HBS is capable of automatically detecting testbench targets. Testbench targets are targets which names:
@@
For example, for the edge detector from Listing 8, the hbs program detects testbench targets presented in Listing 9.
Listing 9: Testbench targets listed for example edge detector core.
[user@host tmp]\$ hbs list-tb
@@ 4.5. Running targets
HBS allows running any target of registered cores. Even if the target itself has nothing to do with the hardware design. For example, running target print from Listing 10 results with the output shown in Listing 11.
Listing 10: Example print target not related to hardware build.
@@
Listing 11: Running print target result.
@@
However, in most cases, the user wants to run a target related to the flow of the set EDA tool. In such a case, instead of manually calling all of the required tool commands, the user can call the hbs::Run procedure in the core target procedure. The hbs::Run procedure has an optional argument accepting the stage after which the tool flow should stop. This is further described in subsection 4.9. After hbs::Run returns, the user can continue processing. For example, the user can run scripts analyzing code coverage or preparing additional reports.
@@ 4.6. Target parameters
As core targets are just Tcl procedures, they can have parameters. Moreover, parameters can have optional default values. Additionally, HBS allows to provide command line arguments to the run target. This is a very convenient feature in build systems. Listing 12 presents a very simplified example.
Listing 12: Example target with an optional parameter.
@@
The core does not build any hardware design. However, the example shows how the build stage can be passed from the command line to an EDA tool. Listing 13 presents output from running the target with different stage parameter values.
Listing 13: Running print target result.
@@
4.7. Target context
An engineer implementing a given core controls the dependencies of the core. However, the core author does not control who will use the core and how. As targets are regular Tcl procedures, there is a need for a mechanism allowing the core author to evaluate the target procedure in the invariant environment. Such a mechanism in HBS is called the target context.
The target context assures that the following variables are not affected by dependee or dependency target execution:
@@ 4.7. Target context
    HDL library,
  2. (2)
    HDL standard,
  3. (3)
    top module name,
  4. (4)
    arguments prefix,
  5. (5)
    arguments suffix,
  6. (6)
    the core path,
  7. (7)
    the target name,
  8. (8)
    the target path.
Listing 14 presents an example of the target context mechanism.
Listing 14: Target context example.
namespace eval pkg {
namespace eval foo {
proc src-foo {} {
hbs::SetLib "lib-foo"
hbs::AddDep pkg::bar::src-bar
puts "foo lib: \$hbs::Lib"
puts "foo core: \$hbs::ThisCorePath"
puts "foo target: \$hbs::ThisTarget"
hbs::Register
namespace eval bar {
proc src-bar {} {
hbs::SetLib "lib-bar"
puts "bar lib: \$hbs::Lib"
puts "bar core: \$hbs::ThisCorePath"
puts "bar target: \$hbs::ThisTarget"
hbs::Register
Listing 15 presents the output from running the pkg::foo:src-foo target. As can be seen, setting library in a target of one core, does not affect library in the target of another core.
Listing 15: Output from target context example.
[user@host tmp] hbs run pkg::foo::src-foo
bar lib: lib-bar
bar core: pkg::bar
bar target: src-bar
foo lib: lib-foo
foo core: pkg::foo
foo target: src-foo
An author of a core can not easily by mistake break execution of another core target procedure. The target context provides the isolation of target procedure evaluation environments.
4.8. Target dependencies
In HBS, targets might depend on other targets instead of cores depending on cores. Such an approach allows for fine-grained control of dependencies.
To declare target dependency, the user must call the hbs::AddDep procedure within the target procedure. The first argument is the dependency path. The remaining arguments are optional and are passed to the dependency procedure as arguments.
To add multiple distinct dependencies, the user must call hbs::AddDep multiple times. The ability to pass custom arguments to dependency was evaluated as much more advantageous than the ability to add multiple dependencies with a single hbs::AddDep call.
The hbs::AddDep internally calls the dependency procedure with the provided arguments. It also tracks dependencies so that generating a dependency graph is possible. Within a single flow, each target procedure can be run at most once with a particular set of arguments. This implies that if multiple target procedures add the same dependency with the same arguments, the dependency procedure is run only once during the first hbs::AddDep call. To enforce some target procedure rerun, the user can always directly call the target. However, enforcing target procedure rerun usually is an alert that a regular Tcl procedure shall be used instead of the core target procedure.
Listing 16 contains an example core definitions for presenting target dependency rules. There are four cores: core-a, core-b, core-c, generator-core. core-a depends on core-b and core-c. core-b depends on core-c. Moreover, cores core-a and core-b depends on the generator-core. However, they use different argument values for generation.
Listing 16: Example of target dependency rules.
namespace eval core-a {
proc target {} {
hbs::AddDep core-b::target
hbs::AddDep core-c::target
hbs::AddDep generator-core::gen a
hbs::AddDep generator-core::gen x
puts "core-a::target"
hbs::Register
namespace eval core-b {
proc target {} {
hbs::AddDep core-c::target
hbs::AddDep generator-core::gen b
hbs::AddDep generator-core::gen x
puts "core-b::target"
hbs::Register
namespace eval core-c {
proc target {} {
puts "core-c::target"
hbs::Register
namespace eval generator-core {
proc gen {arg} {
puts "generator-core::gen \$arg"
hbs::Register
Listing 17 presents output from running target core-a::target. As can be seen, the core-c::target is run only once, even though both core-a::target and core-b::target depend on it. This is because core-c::target has no arguments and can be added as a dependency only once. On the other hand, generator-core::gen is run three times. This is because generator-core::gen is added as a dependency four times, and three times with a distinct argument value.
Listing 17: Running target dependency example result.
[user@host tmp]\$ hbs run core-a::target
core-c::target
generator-core::gen b
generator-core::gen x
core-b::target
generator-core::gen a
core-a::target
4.9. EDA tool flow and stages
The primary function of any hardware build system is to provide the ability to build a design. What the term “build” actually means usually depends on the EDA tool. For some, it is only synthesis; for others, it is synthesis and implementation, and yet for others it might be simulation or just code linting. Each EDA tool is characterized by a distinct flow consisting of different stages. To mimic this behavior, HBS supports the following stages (alphabetical order):
  1. (1)
    analysis - HDL files analysis,
  2. (2)
    bitstream - device bitstream generation,
  3. (3)
    elaboration - design elaboration,
  4. (4)
    implementation - design implementation,
  5. (5)
    project - project creation,
  6. (6)
    simulation - design simulation,
  7. (7)
    synthesis - design synthesis.
Not all stages make sense for all EDA tools. Which stages are present in the given tool flow depends on that tool implementation. Namely, on the implementation of the hbs::<tool>::run procedure. This procedure is called at the end of the hbs::Run procedure execution.
Figure 1 shows the structure of the tool flow for the Vivado project mode. There are four stages: project, synthesis, implementation and bitstream. Each stage is wrapped by custom, user-defined callbacks. The number of callbacks is unlimited. Callbacks in any pre or post stage are executed in the order users add them. The user adds a callback by calling a dedicated HBS Application Programming Interface (API) procedure. For example, to add a post synthesis callback user can call the hbs::AddPostSynthCb procedure. Callbacks can be added with custom argument values.
The post stage callbacks and pre stage callbacks for adjacent stages were added for two reasons.
  1. (1)
    To clearly communicate which stage the callback refers to. For example, configuring implementation settings based on the synthesis results can be done in a post-synthesis callback or pre-implementation callback. However, as the callback modifies the implementation settings, it is probably better to add it using the hbs::AddPreImplCb, than hbs::AddPostSynthCb.
  2. (2)
    To introduce, to some extent, a manageable order of callbacks execution. The pre and post callbacks of a given stage are executed in the order they are added. If some nested dependencies add callbacks for the same pre or post stage, then the order of callbacks execution depends on the order of hbs::AddDep calls. However, if the result of one of the callbacks depends on the result of the other one, then relying on a user to call the hbs::AddDep procedures in proper order is error prone. In such a case, the callback that must be executed as the first one can be added to the post-synthesis callbacks, and the second callback can be added to the pre-implementation callbacks. Such an approach is immune to the order of hbs::AddDep calls.
The user can utilize stage callbacks in any desired way. However, the primary purpose of stage callbacks is to adjust the design build based on the results from a particular stage. For example, the user might want to configure additional implementation settings based on the synthesis results. The user might even terminate the tool flow in a given callback and report an error if certain conditions are not met.
Figure 1. Vivado project mode tool flow.
4.10. EDA tool commands custom arguments
The HBS has a minimal common abstraction layer. The user performs all the actions not covered by the abstraction layer by directly calling EDA tool commands. For example, generating extra timing or clock network reports. However, some of the actions are hidden under the HBS API. For example, adding or analyzing HDL files. EDA tool commands used to perform actions hidden under common API usually have some parameters that are not used by default. Nevertheless, sometimes there is a need to specify additional parameters. In such a case, there are two possible solutions.
The first option is to bypass the HBS API and directly call the underlying EDA tool command. The drawback of this approach is that the user must manually handle the current context. For example, when adding an HDL file, the user must manually specify the library or standard revision. Bypassing HBS API also bypasses the target context.
The second option is to set the underlying command arguments prefix or suffix. This can be achieved with the hbs::SetArgsPrefix and hbs::SetArgsSuffix procedures. The argument prefix is always appended after the command name, and the argument suffix is always appended after all command arguments.
Listing 18 presents Ethernet Management Data Input/Output (MDIO) core definition. The core has one testbench target utilizing the nvc simulator. Nvc report messages occupy multiple lines by default. However, this can be changed by specifying the --messages=compact argument when running the simulation. As running the simulation is the last stage of the nvc flow, the call to hbs::SetArgsPrefix must be wrapped by the call to the hbs::AddPostElabCb procedure.
Listing 18: Ethernet MDIO core definition.
namespace eval vhdl::ethernet {
namespace eval mdio {
proc src {} {
hbs::SetLib "ethernet"
hbs::AddFile mdio.vhd
proc tb {} {
hbs::SetTool "nvc"
hbs::AddPostElabCb hbs::SetArgsPrefix "--messages=compact"
src
hbs::SetLib ""
hbs::AddFile tb/tb-mdio.vhd
hbs::SetTop "tb_mdio"
hbs::Run
hbs::Register
Listing 19 shows commands executed by the hbs to run the simulation. The --messages=compact argument was appended right after the nvc command.
Listing 19: Ethernet MDIO testbench commands.
[user@host vhdl-ethernet]\$ hbs run vhdl::ethernet::mdio::tb
nvc --std=2019 -L. --work=ethernet -a vhdl/vhdl-ethernet/mdio.vhd
nvc --std=2019 -L. --work=work -a vhdl/vhdl-ethernet/tb/tb-mdio.vhd
nvc --std=2019 -L. -e tb_mdio
nvc --messages=compact --std=2019 -L. -r tb_mdio --wave
4.11. HBS API extra symbols
The HBS API consists not only of symbols related to the common EDA abstraction layer. For example, there are extra hbs::Exec and hbs::CoreDir procedures. The first one is a wrapper for the Tcl standard exec procedure. Before calling exec, the hbs::Exec changes the working directory to the directory where the currently evaluated core is defined. When exec returns, the hbs::Exec restores the working directory. The hbs::CoreDir procedure allows the user to get the path of the directory in which the currently evaluated core is defined.
HBS also provides users with the following extra variables:
  1. (1)
    hbs::ThisCorePath - the path of the core which target is currently being run,
  2. (2)
    hbs::ThisTargetPath - the path of the target which is currently being run,
  3. (3)
    hbs::ThisTargetName - the name of the target which is currently being run,
  4. (4)
    hbs::TopCorePath - the path of the top core beign run,
  5. (5)
    hbs::TopTargetPath - the path of the top target being run,
  6. (6)
    hbs::TopTarget - the name of the top target being run,
  7. (7)
    hbs::TopTargetArgs - the list with command line arguments passed to the top target.
4.12. Code generation
Code generation in the hardware design domain is omnipresent, as it significantly speeds up the implementation process. For example, hardware-software co-design system on chip projects usually have some tool automatically generating register files (Werner et al., 2020). Even in pure FPGA designs, it is common to generate descriptions at the register transfer level from some higher-level programming abstraction (Cong et al., 2011; Nane et al., 2016). That is why it is important for any hardware build system to provide as simple mechanism for code generation as possible.
Some existing hardware build systems, for example FuseSoc, do not allow the direct calling of an arbitrary external program in arbitrary places during the build process. Instead, the user has to define so-called generators (O., 2025). Only then can the user call the generator wihtin core definitions. However, such an approach has some drawbacks:
  1. (1)
    The generator call requires an extra layer of indirection. Generators are defined in different places than they are used, which decreases the readability of the description.
@@ 4.12. Code generation
In HBS, there is no formal concept of generator. Anything can be a generator, as generators are just regular Tcl procedures. This means that generators can be target procedures (tracked by dependency system) or core internal Tcl procedures (not tracked by the dependnecy system).
Listing 20 presents an example of calling an external code generator tracked by the dependnecy system. In actual usage, the call to the shell echo command would be replaced with a call to the proper code generator program. Calls to the hbs::AddFile are commented out because no EDA tool was set.
Listing 20: Example of HBS generator tracked by dependency system.
@@
Listing 21 presents how to achieve the same result without tracking the generator as a dependency. This task is even more straightforward, as the user can call an external generator program directly in the target procedure.
Listing 21: Example of HBS generator not tracked by dependency system.
@@
5. Hardware Build System Commands
The implemented hbs program provides the user with 11 commands. Listing 22 presents hbs help message. The help and version commands are not discussed in this section, as their purpose is rather obvious.
Listing 22: HBS help message.
Usage:

### v4

@@ 4.4. Ease of calling arbitrary programs in arbitrary places
The first use case originates from the complexity of designing hardware systems. To reduce manual work, engineers often automatically generate parts of the design. For example, register files are automatically generated from SystemRDL (Accellera SYSTEMS INITIATIVE, 2025). In such a case, there must be a way to call an external program that generates hardware description of the register components.
Existing indirect-abstract hardware build systems allow calling external programs. However, they often require users to define a transitional so-called generator, which increases complexity and reduces legibility. Moreover, they usually do not allow calling external programs in arbitrary places. This is further discussed in subsection 6.6.
The second use case includes support for EDA tools without the Tcl interface. Such tools are often open source simulators, for example, nvc (N., 2025), or verilator (W., 2025). These tools have interfaces similar to the standard software compiler. The user simply calls programs with proper arguments and proper order in a shell. Being able to call arbitrary programs in arbitrary places is required to support these kinds of tools without an extra script or Makefile generation.
4.5. Independence from version control system
@@ 5. Hardware Build System architecture
The HBS official user manual is available in its repository (M., 2025). The purpose of this section is to briefly describe the architecture of HBS so that the reader does not have to study the user manual.
The HBS was implemented to evaluate the limitations of the direct-Tcl approach. Nevertheless, implementing a hardware build system using this approach may be a viable strategy for the following reasons:
@@
    Implementing a hardware build system in Tcl allows the execution of build system code during the EDA tool flow. This, in turn, gives direct access to all EDA tool custom commands. Moreover, these custom commands can be evaluated in arbitrary places. This aspect satisfies requirements 4.1 and 4.4. The complexity of existing indirect-abstract build systems for less common tasks results from executing build system code before the EDA tool flow. The build system only prepares scripts that are later run by EDA tools.
@@
    If the EDA tool provides the Tcl interface, then the Tcl shell is provided by the EDA tool vendor. The shell is installed during the installation of vendor tools. This implies that, in some cases, the build system user does not even have to install additional programs. This aspect satisfies requirement 4.3.
@@
    Executing arbitrary programs in arbitrary places in Tcl is very straightforward. There is a dedicated exec command for invoking subprocesses. If executing a subprocess requires prior dynamic arguments evaluation, the exec command call must be prepended with the eval command. Even in Python, invoking a subprocess is not so simple. This Tcl feature satisfies requirement 4.4.
@@ 5.1. Cores and cores detection
When the user executes hbs, all directories under the working directory are recursively scanned to discover all .hbs files (including symbolic links). Files with the .hbs extension are regular Tcl files that are sourced by the hbs script. However, before sourcing .hbs files, the file list is sorted so that scripts with shorter path depth are sourced first. For example, let us assume the following three .hbs files were found:
@@
Such an approach allows controlling when custom symbols (Tcl variables and procedures) are ready to use. For example, if the user has a custom procedure used in multiple .hbs files, then the user can create separate utils.hbs file containing utility procedures, and place it in the project root directory. Within .hbs files, the user usually defines cores and targets, though they are free to include any valid Tcl code.
@@
To register a core, the user must explicitly call the hbs::Register procedure at the end of the core namespace. Such a mechanism helps distinguish regular Tcl namespaces from those representing core definitions.
@@
    The user can easily stick to the VLNV identifiers if required. This is presented in Listing 3. In this case, the flip-flop core path is vendor::library::flip-flop::1.0.
@@
    The user can define arbitrary deep core paths (limited by the Tcl shell). This is presented in Listing 4. In this case, the core path consists of 7 parts.
@@
    The user can nest namespaces to imitate the structure of libraries and packages. This is presented in Listing 5. Three flip-flop cores are defined in the snippet. Listing 6 presents output for listing flip-flop cores.
Listing 3: Example flip-flop core definition with VLNV core path.
@@
Listing 4: Example flip-flop core definition with 7 core path parts.
@@
Listing 5: An example .hbs file with three flip-flops core definitions
@@
Listing 6: Listing flip-flop cores output.
[user@host tmp]\$ hbs ls-cores
@@ 5.2. Targets and targets detection
The HBS automatically detects targets. Targets are all Tcl procedures defined in the scope of core namespaces (namespaces with a call to the hbs::Register procedure). However, to allow users to define custom utility procedures within cores, procedures with names starting with the floor character (_) are not treated as core targets. Listing 7 presents an example edge detector core definition. The core path is vhdl::simple::edge-detector. The core has three targets: src, tb-sync, tb-comb, and one utility procedure _tb. The _tb procedure was defined to share calls common for testbench targets tb-sync and tb-comb. Moreover, all target procedures are also regular Tcl procedures. Such an approach allows for calling them in arbitrary places. The _tb procedure calls the src procedure because the edge_detector.vhd file is required for running the testbench targets.
All targets are represented by a unique target path. The target path consists of the core path and the target name. For example, the src target of the edge detector has the following path vhdl::simple::edge-detector::src.
Listing 7: Example edge detector core definition.
@@ 5.3. Testbench targets
The HBS is capable of automatically detecting testbench targets. Testbench targets are targets whose names:
@@
For example, for the edge detector from Listing 7, the hbs program detects testbench targets presented in Listing 8.
Listing 8: Testbench targets listed for example edge detector core.
[user@host tmp]\$ hbs ls-tb
@@ 5.4. Running targets
The HBS allows running any target of registered cores, even if the target itself has nothing to do with the hardware design. For example, running target print from Listing 9 results with the output shown in Listing 10.
Listing 9: Example print target not related to hardware build.
@@
Listing 10: Running print target result.
@@
However, in most cases, the user wants to run a target related to the flow of the set EDA tool. In such a case, instead of manually calling all required tool commands, the user can call the hbs::Run procedure from the core target procedure.
@@ 5.5. Target parameters
As core targets are just Tcl procedures, they can have parameters. Moreover, parameters can have optional default values. Additionally, HBS allows providing command-line arguments to the run target. This is a very convenient feature in build systems. Listing 11 presents a very simplified example.
Listing 11: Example target with an optional parameter.
@@
The core does not build any hardware design. However, the example shows how the build stage can be passed from the command line to an EDA tool. Listing 12 presents output from running the target with different stage parameter values.
Listing 12: Running print target result.
@@
5.6. Target dependencies
In HBS, targets depend on other targets rather than on cores. Such an approach allows for fine-grained control of dependencies.
To declare target dependency, the user must call the hbs::AddDep procedure within the target procedure. The first argument is the dependency path. The remaining arguments are optional and are forwarded to the dependency target procedure.
To add multiple distinct dependencies, the user must call hbs::AddDep multiple times. The ability to pass custom arguments to a dependency was evaluated as much more advantageous than the ability to add multiple dependencies with a single hbs::AddDep call.
The hbs::AddDep internally calls the dependency procedure with the provided arguments. It also tracks dependencies, enabling the generation of a dependency graph. Within a single flow, each target procedure can be run at most once with a particular set of arguments. This implies that if multiple target procedures add the same dependency with the same arguments, the dependency procedure is run only once during the first hbs::AddDep call. To enforce a target procedure rerun, the user can always call the target procedure directly. However, enforcing target procedure rerun usually is an alert that a regular Tcl procedure shall be used instead of the core target procedure.
5.7. Code generation
Code generation in the hardware design domain is ubiquitous, as it significantly speeds up implementation. For example, hardware-software co-design system-on-chip projects usually have some tool automatically generating register files (Werner et al., 2020). Even in pure FPGA designs, it is common to generate register-transfer-level descriptions from a higher-level programming abstraction (Cong et al., 2011; Nane et al., 2016). That is why any hardware build system needs to provide as simple mechanism for code generation as possible.
Some existing hardware build systems, for example, FuseSoc, do not allow the direct calling of an arbitrary external program ar arbitrary places during the build process. Instead, the user has to define so-called generators (O., 2025). Only then can the user call the generator within core definitions. However, such an approach has some drawbacks:
@@ 5.7. Code generation
    The generator call requires an extra layer of indirection. Generators are defined in different places than where they are used, which decreases the readability of the description.
@@
In HBS, there is no formal concept of a generator. Anything can be a generator, as generators are just regular Tcl procedures. This means that generators can be target procedures (tracked by the dependency system) or core internal Tcl procedures (not tracked by the dependency system).
Listing 13 presents an example of calling an external code generator tracked by the dependency system. In actual use, the call to the shell’s echo command would be replaced with a call to the proper code-generation program. Calls to the hbs::AddFile are commented out because no EDA tool was set.
Listing 13: Example of HBS generator tracked by dependency system.
@@
Listing 14 presents how to achieve the same result without tracking the generator as a dependency. This task is even more straightforward, as the user can call an external generator program directly in the target procedure.
Listing 14: Example of HBS generator not tracked by dependency system.
@@
5.8. Hardware Build System Commands
The implemented hbs program provides the user with 14 commands. Listing 15 presents hbs help message. Describing particular commands is considered out of scope of the article, as it would not bring any substantive value. The commands are listed solely to present the maturity and versatility of the implemented build system. None of the existing indirect-abstract approach hardware build systems support all features present in HBS. However, some have features missing in HBS, for example, SiliconCompiler, HAMMER, or Hog.
Listing 15: HBS help message.
Usage